OpenStudy (anonymous):

the area of a circular rose garden is 154 meters. what is the radius of the garden? remember a= Pi r^2 use Pi=22/7

Parth (parthkohli):

There's an equation to be formed. \( \color{Black}{\Rightarrow a = \pi r^2, a = 154 \Longrightarrow 154 = {22 \over 7} \times r^2}\) Solve for \(r\).

OpenStudy (anonymous):

πd^2/4

OpenStudy (waleed_imtiaz):

r=7(plus minus)